OpenStudy (anonymous):

Find f(5) for this sequence: f(1) = 2 and f(2) = 3, f(n) = f(1) + f(2) + f(n - 1), for n > 2. f(5) = ______

OpenStudy (kirbykirby):

Try writing what is f(5) first: \[f(5)=f(1)+f(2)+f(5-1)=2+3+f(4) \] What is f(4)? Use the definition again.. \[ f(4)=f(1)+f(2)+f(4-1)=2+3+f(3)\] What is f(3)? Try finding it the same way as above.